OpenStudy (anonymous):

f(x)=tanx2x/x-π/2 if x neqπ/2

OpenStudy (anonymous):

First of all I am quite unsure about what the experssion actually looks like. My interpretation is below. \[f(x) = \frac{\tan(x)*2x}{x}-\frac{\pi}{2} \quad \text{if} \quad x\neq \frac{\pi}{2}\] Secondly I am not sure what the assignment calls for. What are you supposed to do with the experssion?

OpenStudy (anonymous):

\[f(x)=\frac{ \tan 2x }{ x-\frac{ \pi }{ 2 } },x \neq \frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

i think find the limits when \(x\neq \pi/2\)